# QCircuitNet: A Large-Scale Hierarchical Dataset for Quantum Algorithm Design

### v1

@@ Abstract
Quantum computing is an emerging field recognized for the significant speedup it offers over classical computing through quantum algorithms. However, designing and implementing quantum algorithms pose challenges due to the complex nature of quantum mechanics and the necessity for precise control over quantum states. Despite the significant advancements in AI, there has been a lack of datasets specifically tailored for this purpose. In this work, we introduce QCircuitNet, the first benchmark and test dataset designed to evaluate AI’s capability in designing and implementing quantum algorithms in the form of quantum circuit codes. Unlike using AI for writing traditional codes, this task is fundamentally different and significantly more complicated due to highly flexible design space and intricate manipulation of qubits. Our key contributions include:
@@
    A general framework which formulates the key features of quantum algorithm design task for Large Language Models.
@@
    Implementation for a wide range of quantum algorithms from basic primitives to advanced applications, with easy extension to more quantum algorithms.
@@
    Promising potential as a training dataset through primitive fine-tuning results.
We observed several interesting experimental phenomena: fine-tuning does not always outperform few-shot learning, and LLMs tend to exhibit consistent error patterns. QCircuitNet provides a comprehensive benchmark for AI-driven quantum algorithm design, offering advantages in model evaluation and improvement, while also revealing some limitations of LLMs in this domain.
@@ 1 Introduction
Quantum computing is an emerging field in recent decades because algorithms on quantum computers may solve problems significantly faster than their classical counterparts. From the perspective of theoretical computer science, the design of quantum algorithms have been investigated in various research directions - see the survey (Dalzell et al., 2023) and the quantum algorithm zoo (Zoo, 2024). However, the design of quantum algorithms on quantum computers has been completed manually by researchers. This process is notably challenging due to highly flexible design space and extreme demands for a comprehensive understanding of mathematical tools and quantum properties.
@@
For these reasons, quantum computing is often considered to have high professional barriers. As the discipline evolves, we aim to explore more possibilities for algorithm design and implementation in the quantum setting. This is aligned with recent advances among AI for Science, including AlphaFold (Jumper et al., 2021), AlphaGeometry (Trinh et al., 2024), etc. Recently, large language models (LLMs) have also become widely applicable among AI for science approaches (Yang et al., 2024; Zhang et al., 2024; Yu et al., 2024). LLMs represent the best practice of sequential modeling methods at current stage. They have an edge over other models in possessing abundant pre-training knowledge and providing human-friendly interfaces which support human-machine collaboration. Therefore, we gear LLMs for quantum algorithm design.
@@
As far as we know, there has not been any dataset for AI in quantum algorithm design. Existing work combining quantum computing and AI mostly targets at exploiting quantum computing for AI; there are some papers applying AI for quantum computing, but they either consider niche problems (Nakayama et al., 2023; Schatzki et al., 2021) or limited functions (Tang et al., 2023; Fürrutter et al., 2024), not quantum algorithm datasets of general interest (see Section 2). However, unlike classical code generation where abundant data exist, the most challenging aspect for quantum algorithm design is the lack of sufficient data, and hence the difficulty of generalization in training AI models. Therefore, datasets for quantum algorithm design are solicited.
@@
Descriptions of quantum algorithms in natural language could be verbose and vague. Mathematical formulas, while precise and succinct, are difficult to verify automatically. To accommodate with LLMs, we make a change of perspective by formulating quantum algorithms as programming languages. This allows for precise representation of a quantum algorithm, enables automatic verification procedure, and bridges the gap between theoretical design and circuit implementations. Furthermore, meaningful quantum algorithms which can be efficiently implemented have no more than polynomially many gates (Poulin et al., 2011), and thus such formulations have the theoretical benefits allowing for scalable representations.
@@
In this work, We propose QCircuitNet, the first comprehensive, structured dataset for quantum algorithm design. Technically, QCircuitNet has the following key contributions:
  • It formulates the task for Large Language Models (LLMs) with a carefully designed framework encompassing the key features of quantum algorithm design, including problem description, quantum circuit codes, classical post-processing, and verification functions. It maintains the black-box nature of oracles and characterizes query complexity properly.
  • It implements a wide range of quantum algorithms from basic primitives and textbook-level algorithms to advanced applications. We demonstrate the compatibility with complex algorithms through Generalized Simon’s Problem and showcase the easy extension to advanced algorithms.
  • It has automatic validation and verification functions, allowing for iterative evaluation without human inspection. This further enables interactive reasoning which may improve the performance.
  • It showcases the potential as a training dataset through primitive fine-tuning results. As we expand the dataset to include more algorithms and explore novel fine-tuning methods, it will hopefully contribute to interactive quantum algorithm design and implementation significantly.
@@ 2 Related Work
To the best of our knowledge, QCircuitNet is the first dataset tailored specifically for quantum algorithm design. Previous efforts combining quantum computing with AI primarily fall under the category of Quantum Machine Learning (QML), which aims at leveraging the unique properties of quantum systems to enhance machine learning algorithms and achieve improvements over their classical counterparts (Schuld et al., 2015; Biamonte et al., 2017; Ciliberto et al., 2018). Corresponding datasets often focus on encoding classical data into quantum states. For instance, MNISQ (Placidi et al., 2023) is a dataset of quantum circuits representing the original MNIST dataset (LeCun et al., 1998) generated by the AQCE algorithm (Shirakawa et al., 2021). Considering the intrinsic nature of quantum properties, another category of datasets focuses on collecting quantum data to demonstrate quantum advantages since classical machine learning methods can fail to characterize particular patterns of quantum data. For example, Nakayama et al. (2023) created a VQE-generated quantum circuit dataset for classification of variational ansatzes and showed its quantum supremacy. NTangled (Schatzki et al., 2021) further investigated different types of entanglement and composed quantum states with various multipartite entanglement for classification. While these datasets successfully demonstrate the supremacy of quantum computing, they address relatively niche problems whose practical applications are unclear.
@@
This research direction explores the possibility of leveraging AI to facilitate the advancement of quantum computing. QDataSet (Perrier et al., 2022) collects data from simulations of one- and two-qubit systems and targets training classical machine learning algorithms for quantum control, quantum tomography, and noise mitigation. LLM4QPE (Tang et al., 2023) is a large language model style paradigm for predicting quantum system properties with pre-training and fine-tuning workflows. While the paradigm is interesting, the empirical experiments are limited to two downstream tasks: quantum phase classification and correlation prediction. Fürrutter et al. (2024) studied the application of diffusion models (Sohl-Dickstein et al., 2015; Rombach et al., 2022) to quantum circuit synthesis (Saeedi & Markov, 2013; J. et al., 2022). Scalability issues must be addressed to achieve practical and meaningful unitary compilation through this methodology.
@@
The aforementioned works represent meaningful explorations at the intersection of AI and quantum computing. However, none of them considers the task which interests the quantum computing community (from the theoretical side) the most: quantum algorithm design. Our work aims to take the first step in bridging this gap. It is worth noting that several quantum algorithm circuit benchmarks already exist, such as QASMBench (Li et al., 2023), MQTBench (Quetschlich et al., 2023), and VeriQBench (Chen et al., 2022). However, these benchmarks are designed to evaluate the performance of NISQ (Noisy Intermediate-Scale Quantum) (Preskill, 2018) machines or quantum software tools, rather than for training and evaluating AI models. For instance, QASMBench includes a diverse variety of quantum circuits based on OpenQASM representation (Cross et al., 2022), covering quantum circuits with qubit sizes ranging from 2 to 127. However, it fails as a dataset for AI in that it does not capture the design patterns of each algorithm and ignores the post-processing procedure and construction of different oracles, which are crucial to quantum algorithm design. Similar limitations apply to MQTBench and VeriQBench.
3 Preliminaries for Quantum Computing
In this section, we will introduce necessary backgrounds for quantum computing related to this paper. Additional preliminaries can also be found in Appendix B. A more detailed introduction to quantum computing can be found in the standard textbook by Nielsen & Chuang (2000).
Quantum States.
[…]
Quantum Oracles.
$U_{f}|x\rangle|z\rangle=|x\rangle|z\oplus f(x)\rangle,$ (3)
Quantum Gates.
4 QCircuitNet Dataset
@@ 4.1 Task Suite
For the general purpose of quantum algorithm design, we consider three categories of tasks: oracle construction, algorithm design, and random circuit synthesis. These three tasks are crucial for devising and implementing quantum algorithms, with oracle construction serving as the premise for algorithm design, and random circuits serving as a main demonstration for quantum supremacy.
@@ 4.1.1 Task I: Oracle Construction
The construction of such an oracle $U_{f}$ using quantum gates is deeply rooted in the research topic of reversible quantum logic synthesis, which remains a challenge for complex Boolean functions. In this dataset, we mainly focus on the construction of textbook-level oracles: Bernstein-Vazirani Problem (Bernstein & Vazirani, 1993), Deutsch-Jozsa Problem (Deutsch & Jozsa, 1992), Simon’s Problem (Simon, 1997), and Grover’s algorithm for unstructured search (Grover, 1996) (including constructions of both the oracle and the diffusion operator). There is another category of more flexible oracle construction tasks which we refer to as "Problem Encoding". For example, one can apply Grover’s oracle to solving constraint problems such as SAT and triangle finding (Ambainis, 2004). Formulating problem encoding tasks for LLMs slightly differs from quantum logic synthesis, and we refer the readers to Appendix A.2 for more detailed discussion.
@@ 4.1.2 Task II: Quantum Algorithm Design
There are several challenges to address in order to formulate quantum algorithm design task precisely:
  • From a theoretical perspective, the oracle is usually provided as a blackbox gate since the goal of many algorithms is to determine the property of the function $f(x)$ encoded by the oracle $U_{f}$. If the model has access to the gate implementation of the oracle, it can directly deduce the property from the circuit, failing the purpose of designing a quantum algorithm to decode the information. However, for all experimental platforms, a quantum circuit needs to be explicitly constructed to compile and run successfully, i.e., the oracle should be provided with exact gate implementation. Most tutorials and benchmarks (especially those based on Qiskit and OpenQASM) simply merge the circuit implementation of the oracle and the algorithm as a whole for demonstration purposes. When we gear LLMs for quantum algorithm design, how to separate the algorithm circuits from oracle implementation to avoid information leakage is a critical point to consider.
  • A quantum algorithm constitutes not only the quantum circuit, but also its interpretation of measurement results. For example, in Simon’s algorithm, the measurement results $y_{i}$ are not the direct answer $s$ to the problem, but rather satisfies the property $s\cdot y_{i}=0$. Linear equations need to be solved to obtain the final answer. In this case, for a complete algorithm design, the model should also specify the the post-process steps to derive the answer to the original problem.
  • Quantum circuits for the same algorithm vary with different qubit number $n$. Although this is trivial for theoretical design, it needs to be considered when implementing concrete quantum circuits.
In this category, we cover a wide range of quantum algorithms with varying complexity, from fundamental primitives and textbook-level algorithms to advanced applications. For example, we implemented Generalized Simon’s Problem (Ye et al., 2021), a more advanced version of the standard Simon’s problem and an active area of research in recent years (Ye et al., 2021; Wu et al., 2022). The setting is formally stated as follows: given an (unknown) function $f\colon\mathbb{Z}_{p}^{n}\rightarrow X$ where $X$ is a finite set and a $k$ is a positive integer satisfying $k<n$, it is guaranteed that there exists a subgroup $S\leq\mathbb{Z}_{p}^{n}$ of rank $k$ such that for any $x,y\in\mathbb{Z}_{p}^{n},f(x)=f(y)$ iff $x-y\in S$. The goal is to find $S$. Intuitively, the generalized Simon’s problem extends the standard Simon’s problem from binary to $p$-ary bases and from a single secret string to a subgroup of rank $k$. Beyond universal quantum algorithms, we also consider quantum teleportation and quantum key distribution, two widespread protocols in quantum information. We cover their details in Appendix B.
@@ 4.1.3 Task III: Random Circuit Synthesis
The third task we consider is random circuit synthesis. On the one hand, random circuit sampling is the first algorithm for showing quantum supremacy by Google in 2019 (Arute et al., 2019), and is still widely applied to demonstrate the power of quantum algorithms in recent research (Wu et al., 2021; Bluvstein et al., 2024; DeCross et al., 2024). On the other hand, this also naturally enlarges our dataset. To be more specific, we generate the dataset by randomly sampling gates from a Clifford gate set {H, S, CNOT} and a universal set {H, S, T, CNOT} separately. We then simulate the circuits to obtain the final quantum states. In each task, the problem description provides the vector of the final state, and the model is required to generate quantum circuits that reproduce this state using the specified gate set.
@@ 4.2 Dataset Structure
The overall structure of QCircuitNet is illustrated as follows (more details are given in Appendix A):
Figure 1: Structure of QCircuitNet. The components of QCircuitNet are presented in the frame on the top-right. As a showcase, this figure presents the components for Simon’s problem (Simon, 1997), including its problem description in natural language, post-processing function in python code, circuit in a .qasm file, and oracle definition in a .inc file.
@@
As discussed in Section 4.1, a critical consideration in formulating the framework is the dilemma between providing the oracle as a black box for quantum algorithm design and the need for its explicit construction to execute the circuit and interpret the results, making the algorithm design complete. Additionally, model training and reference present challenges, particularly for LLMs in generating complex and precise composite gates and evaluating the results efficiently. To address these obstacles, we highlight the following construction principles, which are specially designed to adapt to these tasks:
  • For algorithm design tasks, as discussed in Section 4.1.2, we provide the oracle as a black-box gate named "Oracle" with the explicit definition in a separate "oracle.inc" library, which is supported by the OpenQASM 3.0 grammar. In this way, we make sure that the model can use the oracle without accessing its underlying function, which solves the problem of isolating oracle definition from the algorithm circuit.
  • For oracle construction tasks, we ask the model to directly output the quantum circuit in QASM format. For algorithm design task, we require both a quantum circuit and a post-processing function to derive the final answer from circuit execution results. Moreover, we ask the model to explicitly set the shots needed to run the circuit itself in order to characterize the query complexity, which is critical in the theoretical analysis of algorithms.
  • For available quantum gates, we provide the definition of some important composite gates not included in the standard QASM gate library in a "customgates.inc". Hierarchical definition for multi-controlled X gate contains 45060 lines for qubit number $n$ = 14 in OpenQASM format, which is impossible for AI models to accurately generate at the time. Providing these as a .inc file guarantees the correctness of OpenQASM’s grammar while avoiding the generation of complicated gates, which is a distraction from the original design task.
  • To verify models’ output automatically without human evaluation, we compose verification functions to validate the syntax of QASM / Qiskit and the functionality of the implemented circuits / codes. Since comprehensive Logic Equivalence Checking (LEC) might be inefficient for the throughput of LLM inference, we perform the verification by directly checking the correctness of output with extensive test cases.
Based on theses principles, we proposed the framework of QCircuitNet. Below is a more detailed explanation for the 7 components of the dataset:
@@
    Generation Code: one general Qiskit (Javadi-Abhari et al., 2024) code to create quantum circuits for oracles or algorithms of different settings, such as distinct secret strings or various qubit numbers. We choose Qiskit as the main experiment platform because it is a general quantum programming software widely used for the complete workflow from creating quantum circuits to transpiling, simulation, and execution on real hardware.
@@
    Algorithm Circuit: a .qasm file storing the quantum circuit for each specific setting. We choose OpenQASM 3.0 (Cross et al., 2022) as the format to store the quantum circuits, because Qiskit, as a python library, can only create quantum circuits at runtime instead of explicitly saving the circuits at gate level. [^1] Although currently the Qiskit APIs for importing and dumping OpenQASM 3.0 files are still in experimental stage, we choose to adopt version 3.0 over 2.0 in that it supports saving parameterized circuits, which allows for extending the framework to variational quantum algorithms (Cerezo et al., 2021).
@@
    Post-Processing Function: this is for Algorithm Design task only, see Section 4.1.2. The function takes a complete quantum circuit as input, uses the Qiskit AerSimulator to execute the circuit, and returns the final answer to the original problem according to the simulation results. For state preparation problems such as creating a GHZ state of $n$ qubits, this function returns the qubit indices of the generated state.
@@
    Oracle / Gate Definition: a .inc file to provide definitions of composite gates or oracles. For oracle construction tasks, this only includes the definition of composite gates required to build the oracle. For algorithm design tasks, we also provide the gate definition of the oracle in this file, which successfully delivers the oracle in a black-box way.
@@
    Verification Function: a function to evaluate whether the implemented oracle / algorithm achieves the desired purpose with grammar validation and test cases verification. If there exist grammar errors, the function returns -1 and provides detailed error message, which can be used as feedback for the LLM to improve through interactive reasoning. If the program can execute successfully, the function returns a score between $[0,1]$ indicating the success rate on test cases. [^2] The verification function explicitly integrates the oracle / gate definition library with output algorithm circuit since Qiskit importer for OpenQASM 3.0 does not support non-standard gate libraries currently.
@@
    Dataset Creation Script: the script to create the dataset from scratch in the format suitable for benchmarking / fine-tuning LLMs. It contains the following functions: 1. generate primitive QASM circuits. 2. extract gate definitions and add include instructions to create algorithm circuit, the direct output of model. 3. validate and verify the correctness of the data points in the dataset. 4. concatenate algorithm circuit with problem description as a json file for the benchmark pipeline.
This structure of QCircuitNet provides a general framework to formulate quantum algorithm design for large language models, with an easy extension to more advanced quantum algorithms.

### v2

# QCircuitBench: A Large-Scale Dataset for Benchmarking Quantum Algorithm Design

@@ Abstract
Quantum computing is an emerging field recognized for the significant speedup it offers over classical computing through quantum algorithms. However, designing and implementing quantum algorithms pose challenges due to the complex nature of quantum mechanics and the necessity for precise control over quantum states. Despite the significant advancements in AI, there has been a lack of datasets specifically tailored for this purpose. In this work, we introduce QCircuitBench, the first benchmark dataset designed to evaluate AI’s capability in designing and implementing quantum algorithms using quantum programming languages. Unlike using AI for writing traditional codes, this task is fundamentally more complicated due to highly flexible design space. Our key contributions include:
@@
    A general framework which formulates the key features of quantum algorithm design for Large Language Models.
@@
    Implementations for quantum algorithms from basic primitives to advanced applications, spanning 3 task suites, 25 algorithms, and 120,290 data points.
@@
    Promising potential as a training dataset through preliminary fine-tuning results.
We observed several interesting experimental phenomena: LLMs tend to exhibit consistent error patterns, and fine-tuning does not always outperform few-shot learning. In all, QCircuitBench is a comprehensive benchmark for LLM-driven quantum algorithm design, and it reveals limitations of LLMs in this domain.
@@ 1 Introduction
Quantum computing is an emerging field in recent decades because algorithms on quantum computers may solve problems significantly faster than their classical counterparts. From the perspective of theoretical computer science, the design of quantum algorithms have been investigated in various research directions - see the survey (dalzell2023quantum) and the quantum algorithm zoo (Zoo). However, the design of quantum algorithms on quantum computers has been completed manually by researchers. This process is notably challenging due to highly flexible design space and extreme demands for a comprehensive understanding of mathematical tools and quantum properties.
@@
For these reasons, quantum computing is often considered to have high professional barriers. As the discipline evolves, we aim to explore more possibilities for algorithm design and implementation in the quantum setting. This is aligned with recent advances in AI for Science, including AlphaFold (jumper2021highly), AlphaGeometry (trinh2024solving), etc. Specifically, large language models (LLMs) have emerged as powerful tools in this domain (yang2024leandojo; zhang2024xiwu; yu2024llasmol; ren2025deepseekprover; lin2025goedel). LLMs represent the best practice of sequential modeling methods at current stage. They have an edge over other models in possessing abundant pre-training knowledge and providing human-friendly interfaces which support human-machine collaboration. Therefore, we employ LLMs as a core component for facilitating quantum algorithm design.
@@
To the best of our knowledge, there is no existing dataset specifically designed for AI-driven quantum algorithm design. Existing work combining quantum computing and AI mostly targets at exploiting quantum computing for AI; there are some papers applying AI for quantum computing, but they either consider niche problems (nakayama2023vqe; schatzki2021entangled) or limited functions (tang2023q; furrutter2024quantum), not quantum algorithm datasets of general interest (see Section 2). Unlike classical code generation where abundant data exist, the most challenging aspect for quantum algorithm design is the lack of sufficient data, and hence the difficulty of generalization in training AI models. Therefore, datasets for quantum algorithm design are solicited.
@@
Choosing an appropriate representation is critical for enabling both precise reasoning and practical usability in quantum algorithm design. Descriptions of quantum algorithms in natural language can be verbose and ambiguous. Mathematical formulas, while precise and compact, are difficult to parse and verify automatically. To accommodate with LLMs, we make a change of perspective by formulating quantum algorithm design as quantum code generation. This allows for precise representation of a quantum algorithm, enables automatic verification procedure, and bridges the gap between theoretical design and circuit implementations. Furthermore, meaningful quantum algorithms which can be efficiently implemented have no more than polynomially many gates (poulin2011quantum), and thus such formulations have the theoretical benefits of allowing for scalable representations.
@@
We propose QCircuitBench, the first comprehensive, structured dataset for quantum algorithm design. Technically, it has the following key contributions:
  • It formulates the task for LLMs with a carefully designed framework encompassing the key features of quantum algorithm design, including problem description, quantum circuit codes, classical post-processing, and verification functions. It maintains the black-box nature of oracles and characterizes query complexity properly.
  • It implements a wide range of quantum algorithms, covering 3 task suites, 25 algorithms, and 120,290 data points. The dataset spans from basic primitives and textbook-level algorithms to advanced applications such as Generalized Simon’s Problem, demonstrating compatibility with complex algorithms and easy extensibility.
  • It has automatic validation and verification functions, enabling iterative, human-free evaluation and supporting interactive reasoning to enhance performance.
  • It showcases the potential as a training dataset through preliminary fine-tuning results. As we expand the dataset to include more algorithms and explore novel fine-tuning methods, it will hopefully contribute to interactive quantum algorithm design and implementation significantly.
@@ 2 Related Work
To the best of our knowledge, QCircuitBench is the first dataset tailored specifically for quantum algorithm design. Previous efforts combining quantum computing with AI primarily fall under the category of Quantum Machine Learning (QML), which aims at leveraging the unique properties of quantum systems to enhance machine learning algorithms and achieve improvements over their classical counterparts (schuld2015introduction; biamonte2017quantum; ciliberto2018quantum). Corresponding datasets often focus on encoding classical data into quantum states. For instance, MNISQ (placidi2023mnisq) is a dataset of quantum circuits representing the original MNIST dataset (lecun1998gradient) generated by the AQCE algorithm (shirakawa2021automatic). Another category of datasets focuses on collecting quantum data to demonstrate quantum advantages since classical machine learning methods can fail to characterize particular patterns of quantum data. nakayama2023vqe created a VQE-generated quantum circuit dataset for classification of variational ansatzes. NTangled (schatzki2021entangled) further investigated different types of entanglement and composed quantum states with various multipartite entanglement for classification. While these datasets successfully demonstrate quantum supremacy, the practical applications of the problem addressed are unclear.
@@
This research direction explores the possibility of leveraging AI to facilitate the advancement of quantum computing. QDataSet (perrier2022qdataset) collects data from simulations of one- and two-qubit systems and targets training classical machine learning algorithms for quantum control, quantum tomography, and noise mitigation. LLM4QPE (tang2023q) is a large language model style paradigm for predicting quantum system properties with pre-training and fine-tuning workflows. While the paradigm is interesting, the empirical experiments are limited to two downstream tasks: quantum phase classification and correlation prediction. furrutter2024quantum studied the application of diffusion models (sohl2015deep; rombach2022high) to quantum circuit synthesis (saeedi2013synthesis; osti_1993220). Scalability issues must be addressed to achieve practical and meaningful unitary compilation through this methodology.
@@
The aforementioned works represent meaningful explorations at the intersection of AI and quantum computing. However, none of them considers the task which interests the quantum computing community (from the theoretical side) the most: quantum algorithm design. Our work aims to take the first step in bridging this gap. It is worth noting that several quantum algorithm circuit benchmarks already exist, such as QASMBench (li2023qasmbench), MQTBench (quetschlich2023mqtbench), and VeriQBench (chen2022veriqbench). However, these benchmarks are designed to evaluate the performance of NISQ (Noisy Intermediate-Scale Quantum) (preskill2018quantum) machines or quantum software tools, rather than for training and evaluating AI models. For instance, QASMBench includes a diverse variety of quantum circuits based on OpenQASM representation (cross2022openqasm), covering quantum circuits with qubit sizes ranging from 2 to 127. However, it fails as a dataset for AI in that it includes only a few entries for each algorithm and ignores the post-processing procedure and construction of different oracles, which are crucial to quantum algorithm design. Similar limitations apply to MQTBench and VeriQBench.
3 QCircuitBench Dataset
@@ 3.1 Task Suite
For the general purpose of quantum algorithm design, we consider three categories of tasks: oracle construction, quantum algorithm design, and random circuit synthesis. These tasks are crucial for devising and implementing quantum algorithms, with oracle construction serving as the premise for algorithm design, and random circuits serving as a main demonstration for quantum supremacy. These task suites encompass 25 algorithms and a total of 120,290 data points with the following distribution:
@@ 3.1.1 Task I: Oracle Construction
This task suite contains 35,872 data points in total, focused on two types of oracle constructions.
To study a Boolean function $f\colon\{0,1\}^{n}\to\{0,1\}^{m}$, we need to gain its access. In quantum computing, the function $f$ is encoded as an oracle $U_{f}$ such that for any $x\in\{0,1\}^{n}$, $z\in\{0,1\}^{m}$, $U_{f}|x\rangle|z\rangle=|x\rangle|z\oplus f(x)\rangle$, where $\oplus$ is the plus modulo 2. The construction of $U_{f}$ using quantum gates is deeply rooted in reversible quantum logic synthesis, which remains a challenge for complex Boolean functions. In this dataset, we mainly focus on the construction of textbook-level oracles: Bernstein-Vazirani Problem (bernstein1993quantum), Deutsch-Jozsa Problem (deutsch1992rapid), Simon’s Problem (simon1997power), and Grover’s algorithm for unstructured search (grover1996fast) (including constructions of both the oracle and the diffusion operator).
There is another category of more flexible oracle construction tasks which we refer to as "Problem Encoding". For example, one can apply Grover’s oracle to solving constraint problems such as SAT and triangle finding (ambainis2004quantum). Formulating problem encoding tasks for LLMs slightly differs from quantum logic synthesis, and we refer the readers to Appendix A.4 for more detailed discussion.
@@ 3.1.2 Task II: Quantum Algorithm Design
In this category, we cover a wide range of quantum algorithms with varying complexity, from fundamental primitives to advanced applications, covering 6,534 data points:
  • Textbook-level algorithms: These range from the Bernstein-Vazirani problem (bernstein1993quantum), Deutsch-Jozsa problem (deutsch1992rapid), Simon’s problem (simon1997power), Grover’s algorithm (grover1996fast), phase estimation (kitaev1995quantum), quantum Fourier transform (coppersmith2002approximate), GHZ state preparation (greenberger2007goingbellstheorem), W state preparation (W2000three), random number generator (herrero2017quantum), swap test (barenco1997stabilization; buhrman2001quantum) to Shor’s algorithm (shor1999polynomial) for factorization, one of the most famous quantum algorithms with superpolynomial speedup.
  • Generalized Simon’s Problem (ye2021query): This is a more advanced version of the standard Simon’s problem and an active area of research in recent years (ye2021query; wu2022simon). The setting is formally stated as follows: given an unknown function $f\colon\mathbb{Z}_{p}^{n}\rightarrow X$, where $X$ is a finite set and a $k$ is a positive integer satisfying $k<n$, it is guaranteed that there exists a subgroup $S\leq\mathbb{Z}_{p}^{n}$ of rank $k$ such that for any $x,y\in\mathbb{Z}_{p}^{n},f(x)=f(y)$ iff $x-y\in S$. The goal is to find $S$. Intuitively, the generalized Simon’s problem extends the standard Simon’s problem from binary to $p$-ary bases and from a single secret string to a subgroup of rank $k$. We conduct experiments for the case $p=3$ (ternary).
  • Variational quantum algorithms (VQAs): Beyond universal quantum algorithms, VQAs including VQE (peruzzo2014variational) for ground-state energy estimation, QAOA (farhi2014quantum) for combinatorial optimization, QAE (romero2017quantum) for quantum state compression, and ENC (LaRose2020robust) for encoding classical data into quantum representations, are potentially implementable on near-term quantum computers. Unlike fixed-circuit quantum algorithms, VQAs require iterative optimization over parameterized circuits, challenging models to generate both circuit structures and effective initialization and optimization procedures.
  • Quantum information protocols: Additionally, we also include quantum information protocols such as quantum teleportation (bennett1993teleporting) and quantum key distribution (bennett2014quantum), which have wide applications in quantum communications, quantum cryptography, etc. See Appendix B for further details.
@@ 3.1.3 Task III: Random Circuit Synthesis
The third task we consider is random circuit synthesis, containing 77,884 data points. On the one hand, random circuit sampling is the first algorithm for showing quantum supremacy by Google (Arute2019), and is still widely applied to demonstrate the power of quantum algorithms in recent research (wu2021strong; bluvstein2024logical; decross2024computational). In this suite, circuits are randomly sampled from a Clifford gate set {H, S, CNOT} and a universal set {H, S, T, CNOT}, and the task is to generate circuits reproducing the specified quantum state.
@@ 3.2 Dataset Structure
The overall structure of QCircuitBench is illustrated in Figure 1 (more details given in Appendix A).
Figure 1: Structure of QCircuitBench. The components of QCircuitBench are presented in the frame on the top-right. As a showcase, this figure presents the components for Simon’s problem (simon1997power), including its problem description in natural language, post-processing function in python code, circuit in a .qasm file, and oracle definition in a .inc file.
Programming Languages and Platforms.
QCircuitBench is designed to generalize across platforms. We provide two functionally equivalent versions, Qiskit (qiskit2024)+OpenQASM (cross2022openqasm) and Cirq (Cirq2025), so the same tasks, oracles, and verification flow are available in either environment. For concreteness, this subsection describes the Qiskit+OpenQASM version. Similar dataset specification applies to Cirq.
@@
Different tasks encounter different challenges. Here we highlight the following construction principles, which are specially designed to adapt to these tasks:
  • Paradox of Oracle Implementation: Quantum algorithms often treat the oracle $U_{f}$ as a black-box, aiming to deduce properties of function $f(x)$ without directly accessing its implementation. However, quantum circuits for real-world platforms need an explicit gate definition to compile and run successfully. To address this challenge, which is often overlooked in tutorials and benchmarks, we provide the oracle as a black-box gate with its explicit definition in a separate "oracle.inc" library. This complies with OpenQASM syntax while ensures the oracle’s functionality is accessible to the model without exposing its internal structure.
  • Classical Processing Specification: A quantum algorithm involves not only quantum circuits but also the classical processing steps to interpret measurement results. For example, in Simon’s algorithm, the model must solve linear equations $s\cdot y_{i}=0$ from measured $y_{i}$. In addition to quantum circuits, we require the model to specify the classical processing function and define the shot count to characterize query complexity, crucial for the theoretical analysis of the algorithm.
  • Custom Quantum Gates: Some composite gates, not part of the standard QASM library, are essential for advanced algorithms. To avoid model distractions, we provide these custom gates, such as multi-controlled X gates (45,060 lines for 14 qubits), in a "customgates.inc" file. These gates are defined hierarchically, allowing the model to use them without the burden of generating complex gate structures.
  • Automated Verification Function: To ensure that model outputs are syntactically correct and functionally valid, we implement automatic verification tools that check OpenQASM syntax and circuit functionality. Instead of performing exhaustive formal verification or equivalence checking, we use a suite of end-to-end test cases to validate the functional correctness of the generated circuits, enabling efficient model evaluation.
Based on theses principles, we proposed the framework of QCircuitBench. Below is a more detailed explanation for the 7 components of the dataset:
@@
    Generation Code: one general Qiskit code to create quantum circuits for oracles or algorithms of different settings, such as distinct secret strings or various qubit numbers. We choose Qiskit as the main experiment platform because it is a general quantum programming software widely used for the complete workflow from creating quantum circuits to transpiling, simulation, and execution on real hardware.
@@
    Algorithm Circuit: a .qasm file storing the quantum circuit for each specific setting. We choose OpenQASM 3.0 as the format to store the quantum circuits, because Qiskit, as a python library, can only create quantum circuits at runtime instead of explicitly saving the circuits at gate level. [^1] Although currently the Qiskit APIs for importing and dumping OpenQASM 3.0 files are still in experimental stage, we choose to adopt version 3.0 over 2.0 in that it supports saving parameterized circuits, which allows for extending the framework to variational quantum algorithms (cerezo2021variational).
@@
    Post-Processing Function: this is for Quantum Algorithm Design task only, see Section 3.1.2. The function takes a complete quantum circuit as input, uses the Qiskit AerSimulator to execute the circuit, and returns the final answer to the original problem according to the simulation results. For state preparation problems such as creating a GHZ state of $n$ qubits, this function returns the qubit indices of the generated state.
@@
    Oracle/Gate Definition: a .inc file to provide definitions of oracles or composite gates. For oracle construction tasks, this only includes the definition of composite gates required to build the oracle. For algorithm design tasks, we also provide the gate definition of the oracle in this file, which successfully delivers the oracle in a black-box way.
@@
    Verification Function: a function to evaluate whether the implemented oracle or algorithm achieves the desired purpose with grammar validation and semantic verification. The function returns two scores: syntax score and semantic score, with details explained in Section 4.1, Evaluation Metrics. If the program fails to run successfully, a detailed error message is provided as the feedback for LLMs to improve through interactive reasoning. [^2] The verification function explicitly integrates the oracle/gate definition library with output algorithm circuit since Qiskit importer for OpenQASM 3.0 does not support non-standard gate libraries currently. All the verification functions were executed by classical simulations in our experiments, but the APIs we implemented are compatible with IBM hardware and can be easily adapted to real quantum computers.
@@
    Dataset Creation Script: the script to create the dataset from scratch in the format suitable for benchmarking or fine-tuning LLMs. It contains the following functions: 1. generate primitive QASM circuits. 2. extract gate definitions and add include instructions to create an algorithm circuit as the direct output. 3. validate and verify the correctness of the data points in the dataset. 4. concatenate the circuit with problem description as a json file for the benchmark pipeline.
This structure of QCircuitBench provides a general framework to formulate quantum algorithm design for large language models, with an easy extension to more advanced quantum algorithms.